A printer has ink that is 8% cobalt blue color and ink that is 18% cobalt blue color. How many ounces of each ink are needed to make 3 gallon (384 ounces) of ink that is 13% cobalt blue color?

Answer by stanbon(75887)

A printer has ink that is 8% cobalt blue color and ink that is 18% cobalt blue color. How many ounces of each ink are needed to make 3 gallon (384 ounces) of ink that is 13% cobalt blue color?
----
Equations:
a + b = 384 ounces
0.08a + 0.18b = 0.13*384
----
Multiply 1st by 8 and 2nd by 100 to get:
----
8a + 8b = 8*384
8a + 18b = 13*384
=====
Subtract and solve for "b":
10b = 5(384)
b = (1/2)384
b = 192 ounces (amt. of 18% solution needed)
a = 192 Ounces (amt of 8% solution needed)
===================
Cheers,
Stan H.